# FAIR Universe HiggsML Uncertainty Challenge Competition

### v2

@@ 1.1 Background and impact
Ten years ago, part of our team co-organised the Higgs Boson Machine Learning Challenge (HiggsML) [1], the most popular Kaggle challenge at the time attracting 1785 teams. This challenge has significantly heightened interest in applying Machine Learning (ML) techniques within High-Energy Physics and, conversely, has exposed physics issues to the ML community. Whereas previously, the most effective methods predominantly relied on boosted decision trees, Deep Learning has since gained prominence (see, e.g., HEP ML living review). While the LHC has not (yet) discovered new physics beyond the Higgs boson, it has accumulated vast data and will continue to accumulate more data well into the next decade. There is a discovery potential in very precise measurements of particle properties, in particular the Higgs boson which can only be studied at the LHC. This demands meticulous measurement of uncertainties and their incorporation into developing sophisticated AI techniques.
High energy physics (HEP) relies on statistical analysis of aggregated observations. Therefore, the interest in uncertainty-aware ML methods in HEP is nearly as old as the application of ML in the field. Advanced efforts that integrate uncertainties into the ML training include approaches that explicitly depend on nuisance parameters [2, 3, 4, 5, 6, 7, 8, 9], that are insensitive to nuisance parameters [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], that use downstream test statistics in the initial training [28, 29, 30, 31, 32, 33, 34, 35, 36], and that use Bayesian neural networks for estimating uncertainties [37, 38, 39, 40]. Many of these topics were covered in a forward-looking review-type article in Ref. [41]. However, these developments all report technique performance on different ad-hoc datasets, so it is difficult to compare their merits. This competition aims at providing a common ground, with a dataset of sufficient complexity, equipped with systematic bias parameterisations, and a metric.
@@ 1.2 Novelty
This entirely new public competition builds on our experience running several competitions in particle physics and broader. These include the original HiggsML data challenge [1], the TrackML Challenges (NeurIPS 2018 competition) [42, 43], the LHC Olympics [44], AutoML/AutoDL [45, 46], and other competitions. Building on the foundation of the original HiggsML challenge, this competition introduces a significant change by using simulated data that includes biases (or systematic effects) in the test dataset. In addition, participants are asked to provide a confidence interval and not just a point estimate. We have developed an innovative metric to assess their performance.
While there have been previous challenges focusing on meta-learning and transfer-learning, such as the NeurIPS 2021 and 2022 meta-learning challenges [47, 48], Unsupervised and Transfer Learning [49], challenges related to bias e.g. Crowd bias challenge [50], and those addressing distribution shifts, like the Shifts challenge series, and CCAI@UNICT 2023 [51], to the best of our knowledge, this is the first challenge that requires participants to handle systematic uncertainty.
@@ 2 Data
We are using a simulated particle physics dataset for this competition to produce data representative of high energy proton collision data collected by the ATLAS experiment [52] at the Large Hadron Collider (LHC) [53]. The dataset is created using two widely used simulation tools, Pythia 8.2 [54] and Delphes 3.5.0 [55]. We have organised the dataset into a tabular format where each row corresponds to a collision event with 28 features, the measurements recorded from a single proton bunch crossing of interest (see Appendix A). The detailed list of features is available in Appendix C, preceded by a brief introduction to special relativity that can be used to understand them better Appendix B. The events are divided into two categories (see Table 1): signal and background. The signal category includes collision events with a Higgs boson decaying into tau pairs (see Figure 1) (one decaying into a lepton, the other one into a hadronic tau), while the background category includes other processes (subcategories) leading to a similar final state.
@@ 4 Metrics
where $\mu_{16,i}$ and $\mu_{84,i}$ are the bounds of the 68.27% CI for each experiment $i$ within the range $[1,N_{\text{test}}]$.
@@
If the confidence interval $[\mu_{16,i},\mu_{84,i}]$ accurately represents the 68.27% quantile, the true value of $\mu$ should lie within this interval in 68.27% of the pseudo-experiments. Consequently, we employ a penalising function $f$ that penalises models that deviate from this 68.27% reference (Figure 3) :
@@
this parameter establishes a tolerance zone around the anticipated value of $0.6827$, accommodating statistical variances inherent in the evaluation process. We opted for an asymmetric penalty function because, within the field of High Energy Physics (HEP), overestimating uncertainty is deemed more acceptable than underestimating it [56, 57]. Hence, coverage exceeding 68.27% incurs a lesser penalty than coverage falling below 68.27%.
@@
$w$ represents the average width of the Confidence Interval, $c$ is the coverage parameter, and $\epsilon=10^{-2}$ is a regularization term to guard against submissions that report unrealistically narrow CIs. Using the $-\ln()$ function ensures that higher scores are awarded to superior submissions and that the score variations remain within a modest range.
@@ 5 Baselines, code, and material provided
inline,color=green!40]Does the baseline method here still correspond to starting kit
– David
@@ 6.1 Overview
The starting kit is a tutorial notebook designed to familiarise participants with the competition protocol, competition data, scoring, and code to prepare their submissions for the competition website. Participants are also given some ready-to-submit submissions in the GitHub repository to get them started with libraries such as PyTorch [58], TensorFlow [59], scikit-learn [60] etc.
@@ Acknowledgements
We are grateful to the US Department of Energy, Office of High Energy Physics, subprogram on Computational High Energy Physics, for sponsoring this research. Discussions seminal to this work were hosted at Institut Pascal at Université Paris-Saclay with the support of the program “Investissements d’avenir” ANR-11-IDEX-0003-01.
@@ Appendix A Proton collisions and detection
This appendix gives a very minimal introduction to what the Challenge data is about (primarily taken from [61]).
@@
We are using the conventional 3D direct reference frame of ATLAS throughout the document (see Figure 7): the $z$ axis points along the horizontal beam line, and the $x$ and $y$ axes are in the transverse plane with the $y$ axis pointing towards the top of the detector. $\theta$ is the polar angle and $\phi$ is the azimuthal angle. Transverse quantities are quantities projected on the $x-y$ plane, or, equivalently, quantities for which the $z$ component is omitted. Instead of the polar angle $\theta$, we often use the pseudorapidity $\eta=-{\rm ln}\ {\rm tan}(\theta/2)$; $\eta=0$ corresponds to a particle in the $x-y$ plane ($\theta=\pi/2$), $\eta=+\infty$ corresponds to a particle traveling along the $z$-axis ($\theta=0$) direction and $\eta=-\infty$ to the opposite direction ($\theta=\pi$). Particles can be identified in the $\eta$ range in $[-2.5,2.5]$. For $|\eta|\in[2.5,5]$, their momentum is still measured but they cannot be identified. Particles with $|\eta|$ beyond 5 escape detection along the beam pipe.
@@ Appendix B Special relativity
This appendix gives a very minimal introduction to special relativity for a better understanding of how the Higgs boson search is performed and what the extracted features mean (taken mainly from [61]).
@@ B.3 Other useful formulas
where $\phi_{A}-\phi_{B}$ is brought back to the $]-\pi,+\pi]$ range. A good intuition for the $R$ separation is that it behaves like the 3D angle in radians between the two particles.
@@ Appendix C The detailed description of the features
  • All azimuthal $\phi$ angles are in radian in the $]-\pi,+\pi]$ range.
@@
  • Features are indicated as “undefined” when it can happen that they are meaningless or cannot be computed; in this case, their value is $-25$, which is outside the normal range of all variables.
@@
The transverse momentum $\sqrt{p_{x}^{2}+p_{y}^{2}}$ of the leading jet, that is the jet with the largest transverse momentum (undefined if PRI_jet_num$\;=0$).
@@
The pseudorapidity $\eta$ of the leading jet (undefined if PRI_jet_num$\;=0$).
@@
The azimuth angle $\phi$ of the leading jet (undefined if PRI_jet_num$\;=0$).
@@
The transverse momentum $\sqrt{p_{x}^{2}+p_{y}^{2}}$ of the sub leading jet, that is, the jet with the second largest transverse momentum (undefined if PRI_jet_num1absent1\;$\leq$ 1≤ 1).
@@
Variables prefixed with DER (for DERived) are quantities computed from the primitive features on the fly from PRImary features (including possible systematics shifts ) [^8] The code to compute DERived features from PRImitive features can be seen at https://github.com/FAIR-Universe/HEP-Challenge/blob/master/ingestion_program/derived_quantities.py. These quantities were selected by the physicists of ATLAS in the reference document [62] either to select regions of interest or as features for the Boosted Decision Trees used in this analysis in order to enhance signal Higgs boson events separation from background events. DERived features were already present in the HiggsML dataset [61] [^9] The notable exception of DER_mass_MMC which was in the HiggsML dataset but is deliberately absent from the Fair-Universe dataset because it was the result of a complex and lengthy Monte-Carlo Markov Chain integration which is not practical to rerun.). The DERived features correspond to feature engineering; an ideal model to be trained on infinite statistics should not need these features. This distinction between primary and derived features (or "low-level" and "high-level" or "raw variables" and "human-assisted variables") is rather standard in the AI for HEP literature, see for example [63, 64]. There is no guarantee that all DERived features are useful for this challenge (they could even be detrimental in the context of systematics). The challenge participant is free to keep these DERived features, remove them altogether, keep a few, or do more feature engineering.
@@
The absolute value of the pseudorapidity separation (Equation 14) between the two jets (undefined if PRI_jet_num1absent1\;$\leq$ 1≤ 1).
@@
The invariant mass (Equation 12) of the two jets (undefined if PRI_jet_num1absent1\;$\leq$ 1≤ 1).
@@
The modulus (Equation 11) of the vector sum of the missing transverse momenta and the transverse momenta of the hadronic tau, the lepton, the leading jet (if PRI_jet_num1absent1\;$\geq$ 1≥ 1) and the subleading jet (if PRI_jet_num$\;=2$) (but not of any additional jets).
@@
The sum of the moduli (Equation 11) of the transverse momenta of the hadronic tau, the lepton, the leading jet (if PRI_jet_num1absent1\;$\geq$ 1≥ 1) and the subleading jet (if PRI_jet_num$\;=2$) and the other jets (if PRI_jet_num$\;>=3$).
@@
where $A=\sin(\phi_{\text{met}}-\phi_{\text{lep}})*\text{sign}(\sin(\phi_{\text{had}}-\phi_{\text{lep}}))$, $B=\sin(\phi_{\text{had}}-\phi_{\text{met}})*\text{sign}(\sin(\phi_{\text{had}}-\phi_{\text{lep}}))$, and $\phi_{met}$, $\phi_{\text{lep}}$, and $\phi_{\text{had}}$ are the azimuthal angles of the missing transverse energy vector, the lepton, and the hadronic tau, respectively. The centrality is $\sqrt{2}$ if the missing transverse energy vector $\vec{E}_{\text{T}}^{\text{miss}}$ is on the bisector of the transverse momenta of the lepton and the hadronic tau. It decreases to 1 if $\vec{E}_{\text{T}}^{\text{miss}}$ is collinear with one of these vectors and it decreases further to $-\sqrt{2}$ when $\vec{E}_{\text{T}}^{\text{miss}}$ is exactly opposite to the bisector. The logic behind this feature is that if the neutrinos are colinear to the lepton and the hadronic tau (which is a good approximation), then the missing transverse energy vector should be between the lepton and the hadronic tau.
@@
The centrality of the pseudorapidity of the lepton w.r.t. the two jets (undefined if PRI_jet_num1absent1\;$\leq$ 1≤ 1)
@@
The feature list and event sample are primarily inspired from [62]. One crucial difference is that the dataset was produced with a more straightforward (leading-order) event generator (Pythia), and the detector effect was simulated with a more straightforward detector simulation (Delphes rather than Geant4 ATLAS Simulation). These simplifications allowed us to provide to participants a large sample allowing the development of sophisticated models while preserving the complexity of the original problem.
@@ D.2 Impact of biases on features
To detail the impact of the systematics, we need to detail first how the 4-momenta from the final state particles can be reconstructed from the PRImary features, following Appendix B. The four parameters ($P_{x}$,$P_{y}$,$P_{z}$,E) of the four-vector of each particle in the final state can be reconstructed from the PRImary features as follows (using the hadronic tau as an example, and reminding that the mass is neglected so that $E=P$),

### v3

@@ 1.1 Background and impact
Ten years ago, part of our team co-organised the Higgs Boson Machine Learning Challenge (HiggsML) pmlr-v42-cowa14 , the most popular Kaggle challenge at the time attracting 1785 teams. This challenge has significantly heightened interest in applying Machine Learning (ML) techniques within High-Energy Physics and, conversely, has exposed physics issues to the ML community. Whereas previously, the most effective methods predominantly relied on boosted decision trees, Deep Learning has since gained prominence (see, e.g., HEP ML living review). While the LHC has not (yet) discovered new physics beyond the Higgs boson, it has accumulated vast data and will continue to accumulate more data well into the next decade. There is a discovery potential in very precise measurements of particle properties, in particular the Higgs boson which can only be studied at the LHC. This demands meticulous measurement of uncertainties and their incorporation into developing sophisticated AI techniques.
High energy physics (HEP) relies on statistical analysis of aggregated observations. Therefore, the interest in uncertainty-aware ML methods in HEP is nearly as old as the application of ML in the field. Advanced efforts that integrate uncertainties into the ML training include approaches that explicitly depend on nuisance parameters Cranmer:2015bka ; Baldi:2016fzo ; Brehmer:2019xox ; Brehmer:2018hga ; Brehmer:2018kdj ; Brehmer:2018eca ; Nachman:2019dol ; Ghosh:2021roe , that are insensitive to nuisance parameters Blance:2019ibf ; Englert:2018cfo ; Louppe:2016ylz ; Dolen:2016kst ; Moult:2017okx ; Stevens:2013dya ; Shimmin:2017mfk ; Bradshaw:2019ipy ; ATL-PHYS-PUB-2018-014 ; DiscoFever ; Wunsch:2019qbo ; Rogozhnikov:2014zea ; 10.1088/2632-2153/ab9023 ; clavijo2020adversarial ; Kasieczka:2020pil ; Kitouni:2020xgb ; Estrade:2019gzk ; Ghosh:2021hrh , that use downstream test statistics in the initial training Wunsch:2020iuh ; Elwood:2020pik ; Xia:2018kgd ; deCastro:2018mgh ; Charnock_2018 ; Alsing:2019dvb ; Simpson:2022suz ; Feichtinger:2021uff ; Layer:2023lwi , and that use Bayesian neural networks for estimating uncertainties Kasieczka:2020vlh ; Bollweg:2019skg ; Araz:2021wqm ; Bellagente:2021yyh . Many of these topics were covered in a forward-looking review-type article in Ref. Chen:2022pzc . However, these developments all report technique performance on different ad-hoc datasets, so it is difficult to compare their merits. This competition aims at providing a common ground, with a dataset of sufficient complexity, equipped with systematic bias parameterisations, and a metric.
@@ 1.2 Novelty
This entirely new public competition builds on our experience running several competitions in particle physics and broader. These include the original HiggsML data challenge pmlr-v42-cowa14 , the TrackML Challenges (NeurIPS 2018 competition) TrackMLAccuracy2019 ; TrackMLThroughput , the LHC Olympics lhc-olympics , AutoML/AutoDL automl ; autodl , and other competitions. Building on the foundation of the original HiggsML challenge, this competition introduces a significant change by using simulated data that includes biases (or systematic effects) in the test dataset. In addition, participants are asked to provide a confidence interval and not just a point estimate. We have developed an innovative metric to assess their performance.
While there have been previous challenges focusing on meta-learning and transfer-learning, such as the NeurIPS 2021 and 2022 meta-learning challenges elbaz2021pmlr ; carriónojeda2022neurips22 , Unsupervised and Transfer Learning guyon2012analysis , challenges related to bias e.g. Crowd bias challenge crowd-bias-challenge , and those addressing distribution shifts, like the Shifts challenge series, and CCAI@UNICT 2023 ccaiunict-2023 , to the best of our knowledge, this is the first challenge that requires participants to handle systematic uncertainty.
@@ 2 Data
We are using a simulated particle physics dataset for this competition to produce data representative of high energy proton collision data collected by the ATLAS experiment 2008atlas at the Large Hadron Collider (LHC) Evans_2008 . The dataset is created using two widely used simulation tools, Pythia 8.2 Sjostrand:2014zea and Delphes 3.5.0 deFavereau:2013fsa . We have organised the dataset into a tabular format where each row corresponds to a collision event with 28 features, the measurements recorded from a single proton bunch crossing of interest (see Appendix A). The detailed list of features is available in Appendix C, preceded by a brief introduction to special relativity that can be used to understand them better Appendix B. The events are divided into two categories (see Table 1): signal and background. The signal category includes collision events with a Higgs boson decaying into tau pairs (see Figure 1) (one decaying into a lepton, the other one into a hadronic tau), while the background category includes other processes (subcategories) leading to a similar final state.
@@ 4 Metrics
where $\mu_{16,i}$ and $\mu_{84,i}$ are the bounds of the 68.27% CI for each experiment $i$ within the range $1,N_{\text{test}}]$.
@@
If the confidence interval $\mu_{16,i},\mu_{84,i}]$ accurately represents the 68.27% quantile, the true value of $\mu$ should lie within this interval in 68.27% of the pseudo-experiments. Consequently, we employ a penalising function $f$ that penalises models that deviate from this 68.27% reference (Figure 3) :
@@
this parameter establishes a tolerance zone around the anticipated value of $0.6827$, accommodating statistical variances inherent in the evaluation process. We opted for an asymmetric penalty function because, within the field of High Energy Physics (HEP), overestimating uncertainty is deemed more acceptable than underestimating it Nachman:2012zf ; Nachman:2014hsa . Hence, coverage exceeding 68.27% incurs a lesser penalty than coverage falling below 68.27%.
@@
$w$ represents the average width of the Confidence Interval, $c$ is the coverage parameter, and $\epsilon=10^{-2}$ is a regularization term to guard against submissions that report unrealistically narrow CIs. Using the $\ln()$ function ensures that higher scores are awarded to superior submissions and that the score variations remain within a modest range.
@@ 6.1 Overview
The starting kit is a tutorial notebook designed to familiarise participants with the competition protocol, competition data, scoring, and code to prepare their submissions for the competition website. Participants are also given some ready-to-submit submissions in the GitHub repository to get them started with libraries such as PyTorch pytorch , TensorFlow tensorflow , scikit-learn scikit-learn etc.
@@ Acknowledgements
We are grateful to the US Department of Energy, Office of High Energy Physics, subprogram on Computational High Energy Physics, for sponsoring this research, as well as to the ANR Chair of Artificial Intelligence HUMANIA (ANR-19-CHIA-0022). Seminal discussions contributing to this work took place at the workshop “Artificial Intelligence and the Uncertainty Challenge in Fundamental Physics,” sponsored by the CNRS AISSAI Center and the DATAIA Institute, and hosted at Institut Pascal at Université Paris-Saclay. The DATAIA Institute and Institut Pascal are respectively funded by the “Investissements d’Avenir” programs ANR-17-CONV-003 and ANR-11-IDEX-0003-01.
@@ Appendix A Proton collisions and detection
This appendix gives a very minimal introduction to what the Challenge data is about (primarily taken from higgsmldoc ).
@@
We are using the conventional 3D direct reference frame of ATLAS throughout the document (see Figure 7): the $z$ axis points along the horizontal beam line, and the $x$ and $y$ axes are in the transverse plane with the $y$ axis pointing towards the top of the detector. $\theta$ is the polar angle and $\phi$ is the azimuthal angle. Transverse quantities are quantities projected on the xyx-y plane, or, equivalently, quantities for which the $z$ component is omitted. Instead of the polar angle $\theta$, we often use the pseudorapidity $\eta=-{\rm ln}\ {\rm tan}(\theta/2)$; $\eta=0$ corresponds to a particle in the xyx-y plane ($\theta=\pi/2$), $\eta=+\infty$ corresponds to a particle traveling along the $z$-axis ($\theta=0$) direction and $\eta=$-$\infty$ to the opposite direction ($\theta=\pi$). Particles can be identified in the $\eta$ range in [2.5,2.5][-2.5,2.5]. For $|\eta|\in[2.5,5]$, their momentum is still measured but they cannot be identified. Particles with $|\eta|$ beyond 5 escape detection along the beam pipe.
@@ Appendix B Special relativity
This appendix gives a very minimal introduction to special relativity for a better understanding of how the Higgs boson search is performed and what the extracted features mean (taken mainly from higgsmldoc ).
@@ B.3 Other useful formulas
where $\phi_{A}-\phi_{B}$ is brought back to the $\pi,+\pi]$ range. A good intuition for the $R$ separation is that it behaves like the 3D angle in radians between the two particles.
@@ Appendix C The detailed description of the features
  • All azimuthal $\phi$ angles are in radian in the $\pi,+\pi]$ range.
@@
  • Features are indicated as “undefined” when it can happen that they are meaningless or cannot be computed; in this case, their value is 25-25, which is outside the normal range of all variables.
@@
The transverse momentum $\sqrt{p_{x}^{2}+p_{y}^{2}}$ of the leading jet, that is the jet with the largest transverse momentum (undefined if PRI_jet_num=0\;=0).
@@
The pseudorapidity $\eta$ of the leading jet (undefined if PRI_jet_num=0\;=0).
@@
The azimuth angle $\phi$ of the leading jet (undefined if PRI_jet_num=0\;=0).
@@
[…]
@@
Variables prefixed with DER (for DERived) are quantities computed from the primitive features on the fly from PRImary features (including possible systematics shifts ) [^8] The code to compute DERived features from PRImitive features can be seen at https://github.com/FAIR-Universe/HEP-Challenge/blob/master/ingestion_program/derived_quantities.py. These quantities were selected by the physicists of ATLAS in the reference document AtlasHiggsInTauTauPaper15 either to select regions of interest or as features for the Boosted Decision Trees used in this analysis in order to enhance signal Higgs boson events separation from background events. DERived features were already present in the HiggsML dataset higgsmldoc  [^9] The notable exception of DER_mass_MMC which was in the HiggsML dataset but is deliberately absent from the Fair-Universe dataset because it was the result of a complex and lengthy Monte-Carlo Markov Chain integration which is not practical to rerun.). The DERived features correspond to feature engineering; an ideal model to be trained on infinite statistics should not need these features. This distinction between primary and derived features (or "low-level" and "high-level" or "raw variables" and "human-assisted variables") is rather standard in the AI for HEP literature, see for example BaldiExotic ; BaldiTauTau . There is no guarantee that all DERived features are useful for this challenge (they could even be detrimental in the context of systematics). The challenge participant is free to keep these DERived features, remove them altogether, keep a few, or do more feature engineering.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
where $A=\sin(\phi_{\text{met}}-\phi_{\text{lep}})*\text{sign}(\sin(\phi_{\text{had}}-\phi_{\text{lep}}))$, $B=\sin(\phi_{\text{had}}-\phi_{\text{met}})*\text{sign}(\sin(\phi_{\text{had}}-\phi_{\text{lep}}))$, and $\phi_{met}$, $\phi_{\text{lep}}$, and $\phi_{\text{had}}$ are the azimuthal angles of the missing transverse energy vector, the lepton, and the hadronic tau, respectively. The centrality is $\sqrt{2}$ if the missing transverse energy vector $\vec{E}_{\text{T}}^{\text{miss}}$ is on the bisector of the transverse momenta of the lepton and the hadronic tau. It decreases to 1 if $\vec{E}_{\text{T}}^{\text{miss}}$ is collinear with one of these vectors and it decreases further to $\sqrt{2}$ when $\vec{E}_{\text{T}}^{\text{miss}}$ is exactly opposite to the bisector. The logic behind this feature is that if the neutrinos are colinear to the lepton and the hadronic tau (which is a good approximation), then the missing transverse energy vector should be between the lepton and the hadronic tau.
@@
[…]
@@
The feature list and event sample are primarily inspired from AtlasHiggsInTauTauPaper15 . One crucial difference is that the dataset was produced with a more straightforward (leading-order) event generator (Pythia), and the detector effect was simulated with a more straightforward detector simulation (Delphes rather than Geant4 ATLAS Simulation). These simplifications allowed us to provide to participants a large sample allowing the development of sophisticated models while preserving the complexity of the original problem.
@@ D.2 Impact of biases on features
To detail the impact of the systematics, we need to detail first how the 4-momenta from the final state particles can be reconstructed from the PRImary features, following Appendix B. The four parameters ($P_{x}$$P_{y}$$P_{z}$,E) of the four-vector of each particle in the final state can be reconstructed from the PRImary features as follows (using the hadronic tau as an example, and reminding that the mass is neglected so that $E=P$),